OpenStudy (anonymous):

if jorge accidenntly drops a pen from a balcony 64 ft above tue ground and the pens distance above ground is h=-16t squared+s .in how many seconds will the pen reach the ground

OpenStudy (anonymous):

i dont understand this h=-16t squared+s

OpenStudy (anonymous):

Yeah -16 to the second power + s

OpenStudy (anonymous):

This is what I get: since the pen is dropped from 64ft then\[h=-16^2+s=64\]solving for s (seconds), I get:\[-16^2+s=64\rightarrow s=64+16^2=320\]it would take 320 seconds for the pen to hit the ground. Hope I helped, lemme know by giving me a medal :)

OpenStudy (anonymous):

Thanks But On My Hw The Answers are either 2,3,4,or16 seconds

OpenStudy (anonymous):

is that a +s or +5?

OpenStudy (anonymous):

S not 5